OpenStudy (anonymous):

Two lines form a linear system. The slope of one line is 4 and it has a y-intercept of (0, -3). The slope of the other line is 4 and it has a y-intercept of (0, 8). How many solutions does this system have?

OpenStudy (anonymous):

can you put this in slope-intercept form

OpenStudy (anonymous):

no....sorry

OpenStudy (anonymous):

\[y-y1=m(x-x1)\]

OpenStudy (anonymous):

is the awnser one

OpenStudy (anonymous):

I'm not going to give you the direct answer ill help you through it though so you can get the answer

OpenStudy (anonymous):

okedokie

OpenStudy (anonymous):

can you substitute in for the variables

OpenStudy (anonymous):

oh sorry for wasting your time I got the awnser

OpenStudy (anonymous):

wait no I don't sorry I got mixed up

OpenStudy (anonymous):

\[y-y1=m(x-x1)\]

OpenStudy (anonymous):

y=4

OpenStudy (anonymous):

can you plug in the numbers please

OpenStudy (anonymous):

m=4 y1=-3 x1=0

OpenStudy (anonymous):

-3-31=4(0)

OpenStudy (anonymous):

is that right